# Alchemy: Amplifying Theorem-Proving Capability through Symbolic Mutation

### v1

@@ Appendix B Limitations
Data Diversity and Quality
We only define two symbolic rules (using two tactics) to synthesize new theorems. The implementation of the synthesis pipeline is over general and utilizes little domain knowledge, which affects the diversity and quality of synthetic data.
The Cost of Synthesizing
Despite the CPU-only nature of our algorithm, the cost of synthesizing remains huge. We believe the overhead can be significantly reduced with a finer implementation and more specialized tools to interact with the Lean theorem prover.
Single-Round v.s. Multi-Round
Theoretically speaking, our algorithms can be iteratively executed by adding the synthesized theorems into seed theorems. Conversely, the synthesized repository is very heavy, which makes it hard to interact with Lean using Leandojo and deploy our algorithm on existing hardware.
Theorem-level or Term-level
Our method synthesizes theorems from top to bottom and introduces additional state-tactic pairs of specific tactics. Synthesizing formal data at the theorem level is not efficient and not consistent with the step-by-step nature of theorem-proving. Ideally, we anticipate that we can synthesize formal data directly at the term level, which aligns with the characteristics of interactive theorem proving.
Up-to-down v.s. Down-to-up
We synthesize theorems in an up-to-down fashion. We construct the new statements first and then retrieve the correct proofs. The up-to-down fashion depends on a specific set of seed theorems, which restricts the diversity of synthetic data. A more fundamental idea is that we can sample some terms in the symbolic space directly, merge them using symbolic manipulations, and then find the corresponding goals for this new theorem. This AlphaGeometry-style idea is hard to implement in Lean and requires a large amount of domain knowledge and engineering endeavors.
Symbolic Synthesis in Conjunction with Other Techniques
Our proposed method demonstrates significant potential for integration with other techniques to enhance the theorem-proving capabilities of LLMs. We posit that theorem synthesis in the symbolic space serves as a valuable complement to prevailing auto-formalization methods. For instance, it may contribute to the expansion of autoformalized datasets. Besides, our approach generates a substantial quantity of new proven statements which can be utilized as a comprehensive database for Retrieval-Augmented Generation (RAG) (Yang et al., 2023; Wang et al., 2024). Our objective is to amalgamate these methodologies to develop a robust theorem prover in the future.
@@ C.1 Overview
As discussed in Section 3, the entire algorithm is composed of four steps. 1) Find invocable theorems for the candidate theorem by executing a specific tactic and retrieving the resulting proof state. 2) Construct new statements, where we parse the resulting proof state and mutate the old statement with the help of AST. 3) Establish the entire proof by inserting a have tactic and integrating it with the old proof to build the whole proof for this new statement. 4) Verify the correctness of generated theorems in Lean theorem prover. In practice, we separately run the time-consuming first step on hundreds of 8-core CPU nodes and unify step 2) and step 3) together to construct the new theorem. Then we will write back synthetic theorems and run “lake build” to verify the generated theorems.
@@ C.2 Find Invocable Theorems
We set different validation criteria for each tactic. For the rw tactic, if the resulting state is a TacticState, we annotate this theorem as invocable. In contrast, for the apply tactic, the resulting state should be “unsolved goals”. Additionally, we filter the resulting invocable theorems to simplify the problem of constructing new theorems. Specifically, we remove the invocable theorems whose next_state contains meta-variables (e.g.,?$a,?m123$) for the rw tactic and unnamed meta-variables (e.g.,?$e12384$) for the apply tactic. Ultimately, we retrieve the invocable theorems for each candidate theorem. One example of invocable theorems is shown in Fig 5.
@@ C.3 Construct New Theorems
Input: candidate statement $s$, invocable theorem $i$
@@
_$\_,next\_state,inst\leftarrow i$triangleright get the next state and instruction
@@
location $l\leftarrow\textsc{parse}(node,i,flag)$triangleright parse AST node and locate the corresponding part that should to be mutated
$m\leftarrow\textsc{construct}(next\_state)$triangleright parse the next proof state and construct the target string
@@
Input: candidate statement $s$, invocable instruction $i$
@@
_$\_,next\_state,inst\leftarrow i$triangleright get the next state and instruction
$M,G\leftarrow\textsc{parse}(next\_state)$triangleright get the set of metavaribales and other subgoals
for $m\in M$ do $\triangleright$ Assigning metavariables
     Add $\textsc{assign}(m,next\_state)$ to $H$
@@
for $g\in G$ do $\triangleright$ Fill the other subgoals depending on meta-varibales
     Add $\textsc{assign}(g,next\_state,M)$ to $H$
@@
location $l\leftarrow\textsc{parse}(node,i)$triangleright parse AST node and locate the old assumption that needs to be mutated
@@
Constructing new statements for apply tactic is more complex than rw. Applying a theorem may introduce some metavariables and new subgoals into the local context for the resulting proof state as shown in Fig 5. We assign values to the metavariables by parsing the next_state and then retrieve all subgoals containing metavariables as new assumptions. For each new assumption, we can extract its name and type from the proof state. To avoid naming conflicts, we define a set of rules to rename the variable according to the naming conversion of Mathlib [^8] https://leanprover-community.github.io/contribute/naming.html. Ultimately, we concatenate all new assumptions and replace the old assumption with them. This procedure is shown in Algorithm 3.
@@ C.4 Verify the Theorems
Our method creates a set of variants for each candidate theorem in Mathlib4. We write the variants back to the original file and execute lake build for verification. We remove the wrong lines for each file by parsing the error message returned by Lean. Then, we will rebuild the repo to ensure the effectiveness of verification. We remove the files that cause errors in the rebuilding process. Specifically, for each 8-core CPU node, we only build one “.lean” file each time to speed up this process and simplify the logic of parsing. The whole experiment runs on 2,048 CPUs ($256\times 8$-core). The code snippets illustrate the procedure for each CPU node. After verifying the correctness of the synthesized theorem, we extract the state-tactic pairs from our augmented Mathlib repository using Leandojo. For rw or apply, it takes three days for a 96-core CPU machine to trace the enlarged repository. In practice, we split the modified lean files into several portions, separately write them into multiple lean repositories, and trace the repos on several 96-core CPU machines.
@@ C.5 Limitations of Synthesis Pipeline
Our synthesis pipeline is mainly based on the advanced Leandojo tool. We use it to interact with Lean, parse abstract syntax trees and trace state-tactic pairs. However, this tool has the following weaknesses. 1) It will generate a significant number of temporary files that consume substantial disk space when initializing a “dojo” environment. The memory-intensive nature of this tool hinders our ability to effectively implement multiprocessing. 2) Moreover, it lacks native support for tracing a local Lean repository, so we must first upload our data to GitHub. 3) We encounter challenges when tracing a repository of a scale significantly larger than that of Mathlib, which makes it hard to do multi-round synthesis. We aspire to enhance the functionality of the Leandojo tool to tackle more demanding scenarios in our forthcoming endeavors.
@@ D.1 Numerical Analysis
The histogram of the number of variants synthesized by each tactic is shown in Figure 6.
@@
For each tactic, we also list the top 20 theorems with the highest number of variants in Figure 7.
@@ D.2 Examples
Due to the large volume of synthetic data, it is challenging to display all the data in the appendix. We only display a subset of demo theorems for reference. The proof lengths of these theorems range from 1 to 3 lines. To explore further details, please examine our dataset. The synthesized theorems of rw tactic are displayed in Fig 8.
@@
The synthesized theorems of apply are displayed in Fig 9.
@@ D.3.1 Examples of Training Data
As shown in Fig 10, we synthesize a series of variants for each candidate theorem by employing different tactic instructions to mutate existing theorems. We simply combine these additional theorems with the original theorems in Mathlib and train LLMs on this augmented corpus. In addition to synthesizing variants for each candidate theorem, symbolic manipulations to construct new theorems also introduce some new state-tactic pairs. What should be noted is that the state-tactic pairs are extracted by Leandojo rather than manually designed symbolic rules. We have not performed any post-processing on the extracted state-tactic pairs. We group the extracted theorems by the employed tactics (rw, apply, have). The examples of rw and apply are shown in Fig 11. The examples of have are shown in Fig 12.
@@ D.3.4 Influence of the Quantity of SFT Dataset
We assess the impact of varying quantities of additional state-tactics pairs for each tactic under several conditions. 1) Mathlib-train with no additional data points; 2) Downsampling with a ratio of 0.25, resulting in 7.5k additional data points; 3) Downsampling with a ratio of 0.5, resulting in 15k additional data points; 4) Our setting with a deduplication threshold of 1, resulting in 30k additional data points; 5) Deduplication with a threshold of 50, resulting in 500k additional data points; and 6) No deduplication, resulting in 3M additional data points. We fine-tune Llama-3-8b on these different mixtures of data and evaluate their performance on random split of Leandojo Benchmark. The experimental results are shown in Fig 4, demonstrating that our setting achieves a relatively optimal balance between overhead and performance.
@@ E.2.1 Preference in Used Tactics
To see the preference for the tactics used to prove competition-level problems, we perform a comprehensive analysis of the theorems proved by different LLMs. Specifically, we fine-tune different LLMs with the random train-split of Leandojo benchmark and gather all theorems proved by these models. The collection of these models proves 100 theorems out of 244 theorems (41%) on the test split of miniF2F benchmark. The average length of the proofs generated by these models is 1.38. And the distribution of these proved theorems is shown in Fig 14. We have the following observations. 1) About half of the theorems in the miniF2F test split can be proven with only 1-2 line proofs. 2) Most of the theorems are proved with advanced and automatic tactics in Lean (e.g., norm_num, linarith, omega, simp, etc.). We assume that these tactics play an important role in the theorem-proving ability of LLMs to prove competition-level problems. From the above observations, we assume that synthesizing advanced tactic data points rather than basic data points featuring rw and apply is promising to improve the performance of proving competition-level problems.

### v2

@@ Appendix B Limitations
Data Diversity and Quality. We only define two symbolic rules (using two tactics) to synthesize new theorems. The implementation of the synthesis pipeline is over general and utilizes little domain knowledge, which affects the diversity and quality of synthetic data.
The Cost of Synthesizing. Despite the CPU-only nature of our algorithm, the cost of synthesizing remains huge. We believe the overhead can be significantly reduced with a finer implementation and more specialized tools to interact with the Lean theorem prover.
Single-Round v.s. Multi-Round. Theoretically speaking, our algorithms can be iteratively executed by adding the synthesized theorems into seed theorems. Conversely, the synthesized repository is very heavy, which makes it hard to interact with Lean using Leandojo and deploy our algorithm on existing hardware.
Theorem-level or Term-level. Our method synthesizes theorems from top to bottom and introduces additional state-tactic pairs of specific tactics. Synthesizing formal data at the theorem level is not efficient and not consistent with the step-by-step nature of theorem-proving. Ideally, we anticipate that we can synthesize formal data directly at the term level, which aligns with the characteristics of interactive theorem proving.
Up-to-down v.s. Down-to-up. We synthesize theorems in an up-to-down fashion. We construct the new statements first and then retrieve the correct proofs. The up-to-down fashion depends on a specific set of seed theorems, which restricts the diversity of synthetic data. A more fundamental idea is that we can sample some terms in the symbolic space directly, merge them using symbolic manipulations, and then find the corresponding goals for this new theorem. This AlphaGeometry-style idea is hard to implement in Lean and requires a large amount of domain knowledge and engineering endeavors.
Symbolic Synthesis in Conjunction with Other Techniques. Our proposed method demonstrates significant potential for integration with other techniques to enhance the theorem-proving capabilities of LLMs. We posit that theorem synthesis in the symbolic space serves as a valuable complement to prevailing auto-formalization methods. For instance, it may contribute to the expansion of autoformalized datasets. Besides, our approach generates a substantial quantity of new proven statements which can be utilized as a comprehensive database for Retrieval-Augmented Generation (RAG) (Yang et al., 2023; Wang et al., 2024). Our objective is to amalgamate these methodologies to develop a robust theorem prover in the future.
@@ C.1 Overview
As discussed in Section 3, the entire algorithm is composed of four steps: 1) Find invocable theorems for the candidate theorem by executing a specific tactic and retrieving the resulting proof state; 2) Construct new statements, where we parse the resulting proof state and mutate the old statement with the help of AST; 3) Establish the entire proof by inserting a have tactic and integrating it with the old proof to build the whole proof for this new statement; 4) Verify the correctness of generated theorems in Lean theorem prover. In practice, we separately run the time-consuming first step on hundreds of 8-core CPU nodes and unify step 2) and step 3) together to construct the new theorem. Then we will write back synthetic theorems and run “lake build” to verify the generated theorems.
@@ C.2 Find Invocable Theorems
We set different validation criteria for each tactic. For the rw tactic, if the resulting state is a TacticState, we annotate this theorem as invocable. In contrast, for the apply tactic, the resulting state should be “unsolved goals”. Additionally, we filter the resulting invocable theorems to simplify the problem of constructing new theorems. Specifically, we remove the invocable theorems whose next_state contains meta-variables (e.g.,?$a,?m123$) for the rw tactic and unnamed meta-variables (e.g.,?$e12384$) for the apply tactic. Ultimately, we retrieve the invocable theorems for each candidate theorem. One example of invocable theorems is shown in Fig.5.
@@ C.3 Construct New Theorems
Input: candidate statement $s$, invocable theorem $t_{inv}$
@@
_$\_,next\_state,inst\leftarrow t_{inv}$triangleright get the next state and instruction
@@
location $l\leftarrow\textsc{parse}(node,t_{inv},flag)$triangleright parse AST node and locate the corresponding part that should to be mutated
mutant $m\leftarrow\textsc{construct}(next\_state)$triangleright parse the next proof state and construct the target string
@@
Input: candidate statement $s$, invocable theorem $t_{inv}$
@@
_$\_,next\_state,inst\leftarrow t_{inv}$triangleright get the next state and instruction
$Metavs,Goals\leftarrow\textsc{parse}(next\_state)$triangleright get the set of metavaribales and other subgoals
for $metav\in Metavs$ do $\triangleright$ Assigning metavariables
     Add $\textsc{assign}(metav,next\_state)$ to $H$
@@
for $goal\in Goals$ do $\triangleright$ Fill the other subgoals depending on meta-varibales
     Add $\textsc{assign}(goal,next\_state,Metavs)$ to $H$
@@
location $l\leftarrow\textsc{parse}(node,t_{inv})$triangleright parse AST node and locate the old assumption that needs to be mutated
@@
Constructing new statements for apply tactic is more complex than rw. Applying a theorem may introduce some metavariables and new subgoals into the local context for the resulting proof state as shown in Fig.5. We assign values to the metavariables by parsing the next_state and then retrieve all subgoals containing metavariables as new assumptions. For each new assumption, we can extract its name and type from the proof state. To avoid naming conflicts, we define a set of rules to rename the variable according to the naming conversion of Mathlib [^10] https://leanprover-community.github.io/contribute/naming.html. Ultimately, we concatenate all new assumptions and replace the old assumption with them. This procedure is shown in Algorithm 3.
@@ C.4 Verify the Theorems
Our method creates a set of variants for each candidate theorem in Mathlib. We write the variants back to the original file and execute lake build for verification. We remove the wrong lines for each file by parsing the error message returned by Lean. Then, we will rebuild the repo to ensure the effectiveness of verification. We remove the files that cause errors in the rebuilding process. Specifically, for each 8-core CPU node, we only build one “.lean” file each time to speed up this process and simplify the logic of parsing. The whole experiment runs on 2,048 CPUs ($256\times 8$-core). The code snippets illustrate the procedure for each CPU node. After verifying the correctness of the synthesized theorem, we extract the state-tactic pairs from our augmented Mathlib repository using Leandojo. For rw or apply, it takes three days for a 96-core CPU machine to trace the enlarged repository. In practice, we split the modified lean files into several portions, separately write them into multiple lean repositories, and trace the repos on several 96-core CPU machines.
@@ C.5 Limitations of Synthesis Pipeline
Our synthesis pipeline is mainly based on the advanced Leandojo tool. We use it to interact with Lean, parse abstract syntax trees and trace state-tactic pairs. However, this tool has the following weaknesses: 1) It will generate a significant number of temporary files that consume substantial disk space when initializing a “dojo” environment. The memory-intensive nature of this tool hinders our ability to effectively implement multiprocessing; 2) Moreover, it lacks native support for tracing a local Lean repository, so we must first upload our data to GitHub; 3) We encounter challenges when tracing a repository of a scale significantly larger than that of Mathlib, which makes it hard to do multi-round synthesis. We aspire to enhance the functionality of the Leandojo tool to tackle more demanding scenarios in our forthcoming endeavors.
@@ D.1 Numerical Analysis
The histogram of the number of variants synthesized by each tactic is shown in Fig.6.
@@
For each tactic, we also list the top 20 theorems with the highest number of variants in Fig.7.
@@ D.2 Examples
Due to the large volume of synthetic data, it is challenging to display all the data in the appendix. We only display a subset of demo theorems for reference. The proof lengths of these theorems range from 1 to 3 lines. The synthesized theorems of rw tactic are displayed in Fig.8.
@@
The synthesized theorems of apply are displayed in Fig.9.
@@ D.3.1 Examples of Training Data
As shown in Fig.10, we synthesize a series of variants for each candidate theorem by employing different tactic instructions to mutate existing theorems. We simply combine these additional theorems with the original theorems in Mathlib and train LLMs on this augmented corpus. In addition to synthesizing variants for each candidate theorem, symbolic manipulations to construct new theorems also introduce some new state-tactic pairs. What should be noted is that the state-tactic pairs are extracted by Leandojo rather than manually designed symbolic rules. We have not performed any post-processing on the extracted state-tactic pairs. We group the extracted theorems by the employed tactics (rw, apply, have). The examples of rw and apply are shown in Fig.11. The examples of have are shown in Fig.12.
@@ D.3.4 Influence of the Quantity of SFT Dataset
We assess the impact of varying quantities of additional state-tactics pairs for each tactic under several conditions: 1) Mathlib-train with no additional data points; 2) Downsampling with a ratio of 0.25, resulting in 7.5k additional data points; 3) Downsampling with a ratio of 0.5, resulting in 15k additional data points; 4) Our setting with a deduplication threshold of 1, resulting in 30k additional data points; 5) Deduplication with a threshold of 50, resulting in 500k additional data points; and 6) No deduplication, resulting in 3M additional data points. We fine-tune Llama-3-8b on these different mixtures of data and evaluate their performance on random split of Leandojo Benchmark. The experimental results are shown in Fig.4, demonstrating that our setting achieves a relatively optimal balance between overhead and performance.
@@ E.2.1 Preference in Used Tactics
To see the preference for the tactics used to prove competition-level problems, we perform a comprehensive analysis of the theorems proved by different LLMs. Specifically, we fine-tune different LLMs with the random train-split of Leandojo benchmark and gather all theorems proved by these models. The collection of these models proves 100 theorems out of 244 theorems (41%) on the test split of miniF2F benchmark. The average length of the proofs generated by these models is 1.38. And the distribution of these proved theorems is shown in Fig.14. We have the following observations: 1) About half of the theorems in the miniF2F test split can be proven with only 1-2 line proofs; 2) Most of the theorems are proved with advanced and automatic tactics in Lean (e.g., norm_num, linarith, omega, simp, etc.). We assume that these tactics play an important role in the theorem-proving ability of LLMs to prove competition-level problems. From the above observations, we assume that synthesizing advanced tactic data points rather than basic data points featuring rw and apply is promising to improve the performance of proving competition-level problems.